OpenStudy (anonymous):

Which is a solution to the equation 4x + y = –6? A. (–2, –2) B. (2, –2) C. (–2, 2) D. (2, –4) @IMStuck @hartnn @texaschic

OpenStudy (muzzack):

lets do it together

OpenStudy (muzzack):

first we isolate y

OpenStudy (muzzack):

so subtract 4x to both sides

OpenStudy (muzzack):

\[y=-4x-6\]

OpenStudy (muzzack):

now that you do that replace all of the (x,y) and pplug it in the equation and see if it is true or false

OpenStudy (muzzack):

so lets do the first choice

OpenStudy (muzzack):

\[-2=-4(-2)-6\]

OpenStudy (muzzack):

is this true, to find out if it is true simplify it

OpenStudy (anonymous):

nowoi am confused

OpenStudy (muzzack):

-2=8-6 -2=2 false

OpenStudy (muzzack):

now u try it :) hope i helped

OpenStudy (muzzack):

@wakeboarder13 u see the first equation, i made, plug in the (x,y) in to the equation to see if it does equal the same for ex. 5=5

OpenStudy (anonymous):

what would it be

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@Muzzack

OpenStudy (imstuck):

Take each coordinate, one at a time, and fill in its x value and its y value into your equation. Taking the first one, (-2, -2), you get 4(-2) + (-2) = -6. Do the math then and find out if this coordinate makes the equation true. In other words, does sticking in the -2 for x and the -2 for y make the equation = -6.

OpenStudy (anonymous):

so what would it be

OpenStudy (imstuck):

For the first coordinate 4(-2) + (-2) = -6, the result is -10, so the first coordinate, A, is not your answer. Try the second one 4(2)+ (-2) = -6. Does this coordinate make the equation = -6? No this one comes out to a positive 6 and positive 6 does not equal -6. Try the third one, C.

OpenStudy (imstuck):

4(-2)+2 = -6. 4(-2) is -8. -8 + 2 = -6, so C is your answer. Do you understand that?

OpenStudy (anonymous):

its c?

OpenStudy (muzzack):

yess